Seatlle Star blends whole bean coffee worth $0.93 per pound with half bean coffee worth $1.20 per pound to get 30 pounds of a coffee blend worth $1.02 per pound. How many pounds of each type of coffee does she use?
Answer by CharStar(110)

Whole Bean Coffee - $0.93
Amount in lbs = x
Total of Each - 0.93x
Half Bean Coffee - $1.20
Amount in lbs = 30-x
Total Cost of each = $1.20(30-x)
Mixture =$1.02
Amount in lbs = 30
Total Cost of each = $1.02(30)
$0.93x + 1.20(30-x) = $102.(30)
$0.93x + 36.00 - 1.20x = 30.60
-36.00 from both sides and Add the x
$0.27 x = 5.40
x=20
Check your answer
$0.93(20)=$18.60
$1.20(30-20) = $12.00
$18.60 + $12.00 = $36.00
1.02(30) = $36.00
Hope that helps you.